(7) Orthogonally diagonalize the matrix (4 12). 1
Answer #1

4 I orthogonally diagonalize the matrise sols The characteristic equation of the giren is 4- -2 = - 1-> matrise (4-3) = 0 2 (Now Salve the matrix ecuation [: 7] (21-6:] a-k veco a, =%dz let dat -[**]8374 so Pigen rector corresponding Vi = yt to eiLet ♡=t, then ,=-zt V2 = (129-130) + 72=5 is V2 = so, the ergen Vector Corresponding to eigen value [] The diagonal matrise (